### v1

@@ Abstract
By simply applying the Local Potential Approximation (LPA) on the Polchinski’s Exact Renormalization Group (ERG) flow equation for single Bosonic and spinless Fermionic fields, and initially considering only the coarse-graining (blocking) aspect of Wilson’s Renormalization Group program. Within the LPA limit the Polchinski’s ERG flow equation simplifies into a heat differential equation for both Bosonic and Fermionic fields. Solving the differential equations leads to logarithmic interactions (logarithmic vertex function) in both Bosonic and Fermionic fields at their fixed points.
1 Introduction.
Within the Framework of Exact Renormalization Group (ERG) methods [1, 2, 3, 4], with a particular focus on Polchinski’s Exact Renormalization Group Equation (ERGE) [4], which is the Wilson’s Exact Renormalization Group (ERG) [1] derived in the field theory context. We study a system described by a general effective action consisting of single Bosonic and spinless Fermionic field modes including their interactions. Working in Euclidean momentum space where the theory is regularized by employing the cutoff functions of generic form $K_{UV}(p,\Lambda)$ in both fields, whereby the cutoff functions suppress field modes with momentum $p$ that are above the physical UV (ultra-violet) cutoff scale $\Lambda$. The Polchinski’s ERGE emanate from requiring that as the scale $\Lambda$ is lowered by integrating (course-graining) high momentum modes, the effective action (vertex functions) of the theory should be varied in such a way that the partition function of the theory remains unchanged. The variation of the effective action due to lowering of the physical cutoff scale is governed by the Renormalization Group (RG) equation, which is an exact differential equation subsequently giving rise to ERG methods. ERG methods in field theory are exact non-perturbative formulations of RG flow physics encoding it as a differential variation of physical scale and the effective action. They describe the evolution of the renormalized effective action along the RG flow trajectory in terms of differential equations. From the theory setup, these equations arise from the decimation (integration) of fields modes momenta, followed by lowering of the momentum cutoff and then rescaling the action parameters while enforcing the physical observables to remain invariant by requiring the partition function to remain invariant. Usually the RG flow has two physical cutoff scales, one is a high energy scale denoted as the UV cutoff scale, and the second one is a low energy scale denoted as an IR (infrared) cutoff scale, and the RG flow will normally be a flow from the UV (high energy) limit towards the IR (low energy) limit. A generic RG flow equation has the form
@@
Generally solving such an equation specially in the ERGE context can be challenging, since a handful of ERG’s problems end up as integro-differential equations [1, 2, 3, 4], with Polchinski’s ERG flow equation being in this category. Therefore finding an exact non-perturbative general solution from these kind of equations generally it is a difficult tasks, in practice one ends up with an infinite tower of equations and has to always truncate the equations by performing certain approximations. Several non-perturbative approximation solutions and methods are possible, such as derivative expansion, which involves expanding the equations in powers of momenta and this has shown to yield promising results for calculating critical exponents and renormalization group flows equations in $2\leq d\leq 4$ scalar fields theories [5, 6, 7, 8]. The simplest non-perturbative approximation from such power momentum expansion is obtained from taking the leading term (zeroth order term of the expansion), which is the Local Potential Approximation (LPA) expansion. Here we will make use of such power momentum expansion, restrict the computations within the LPA limit. As a start we avoid the consideration of next to leading order terms and other higher order terms in momentum or derivative expansion since these terms are known to yield physical solutions (physical observables) which are scheme dependent (cutoff scale dependent) [8, 9, 10], which makes it difficult to compute physical observables such as the critical exponents ambiguously. Whilst the advantage with the leading order solutions is that, the LPA physical observables solutions are scheme independent and can be computed unambiguously.
It is known and alluded [11] that the Polchinski’s ERG flow equation can be written as a functional heat equation. However, presently it appears there is lack of visible available work towards unpacking this alluded notion at a basic level. Clearly the notion that the Polchinski’s ERG flow equation can be represented as a functional heat equation hasn’t received much attention, and consequently not much work has been done to unpack and elucidate this notion with simple theoretical approaches. Therefore this paper aims to address this, by providing a concrete connection or relation between the Polchinski’s ERG flow equation and the well known standard heat equation within the LPA limit. Utilizing this notion, in one of our main results in section 3 at equation (40), we show that at low energy after relaxing the rescaling steps of Wilson’s Renormalization Group program and only implement the coarse-graining step, subsequently the Polchinski’s ERG flow equation within the LPA limit simplifies into a standard heat equation,
@@
where $U(z,t)=e^{-V(z,t)}$, $t$ is a dimensionless parameter, $z$ represent Bosonic fields and $V(z,t)$ is the vertex or interaction function. We will explain more about the results and how they are obtained in later sections of the paper. It is worth mentioning that the results are interesting due to their simplistic nature of relating the Polchinski’s ERG flow equation within the LPA to a heat equation, this can potentially open a new front to generally study ERG equations through the available technology and language already well developed and established in studying heat equations. Subsequently the solution to this standard heat equation at the fixed point, which is related to the steady state solution in the language of heat equations, we find that the fixed point solution is governed by logarithmic interactions,
@@
where $V_{*}$ is the interaction function at the fixed point obtained at $t=\frac{z^{2}}{2\lambda}$. The solution is interesting since fixed points normally signals a Conformal Field Theory (CFT), therefore at this low energy limit we have established a logarithmic interaction CFT. Contrary to perturbing the solution in the vicinity of the Gaussian fixed point [12, 13, 14, 5], here we perturb the solution in the vicinity of the logarithmic fixed point obtained above. For the perturbation recipe we use a small parameter to linearize the Polchinski’s ERG flow equation in the vicinity of the logarithmic fixed point, but here we implement both rescaling and coarse-graining steps on the Polchinski’s ERG flow equation. This leads to an infinite set tower of Hermite differentials at each expansion order in parameter $t$, with the Hermite polynomials being solutions to these differential equations. Treating the Fermionic fields in similar fashion, in section 4 at equation (77) we obtain the following analogue of a heat equation in Grassmann variables,
@@
The paper is organised as follows. In section 2 we discuss the derivation of Polchinski’s ERG flow equation for a system containing both the single Bosonic and spinless Fermionic fields, and only considering the coarse-graining step of Wislon’s Renormalization Group program. In section 3 we additionally consider the rescaling step of Wilson’s Renormalization Group program on the Polschinki’s ERG equation, we apply the rescaling step and take the LPA limit of the Polchinski’s ERG flow equation, then the differential equation emanating from Polchinki’s ERG flow equation at the LPA limit is solved, followed by finding the eigenvalues, eigenfunctions and the critical exponents for the Bosonic fields sector. Then section 4 deals with solving the Polchinski’s ERG flow equation within the LPA limit in the Fermionic fields sector.
2 Derivation of Polchinski’s ERGE for Bosonic and Fermionic Fields.
@@
$-\frac{S_{int}[\phi,\Lambda]}{\partial\Lambda}=\frac{1}{2}\int\frac{d^{d}p}{(2\pi)^{d}}\frac{\partial G_{B}\big{(}\frac{p}{\Lambda}\big{)}}{\partial\Lambda}\bigg{(}\frac{\delta^{2}S_{int}[\phi,\Lambda]}{\delta\phi(p)\delta\phi(-p)}-\frac{\delta S_{int}[\phi,\Lambda]}{\delta\phi(p)}\frac{\delta S_{int}[\phi,\Lambda]}{\delta\phi(-p)}\bigg{)}$ (14)
$-\Lambda\frac{\partial S_{int}[\phi,\Lambda]}{\partial\Lambda}=\frac{1}{2}\int\frac{d^{d}p}{(2\pi)^{d}}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta^{2}S_{int}[\phi,\Lambda]}{\delta\phi(p)\delta\phi(-p)}-\frac{\delta S_{int}[\phi,\Lambda]}{\delta\phi(p)}\frac{\delta S_{int}[\phi,\Lambda]}{\delta\phi(-p)}\bigg{)},$
@@
$e^{-W_{\Lambda}^{\Lambda_{0}}[J]}=\int D\psi D\bar{\psi}D\phi e^{-\frac{1}{2}\int_{p}\phi(p)G_{B}^{-1}(\frac{p}{\Lambda})\phi(-p)-\int_{p}\bar{\psi}(p)G_{F}^{-1}(\frac{p}{\Lambda})\psi(-p)+\int_{p}J(p)\phi(-p)+\int_{p}\bar{\eta}(p)\psi(-p)+\int_{p}\eta(p)\bar{\psi}(-p)-S_{int}^{\Lambda_{0}}[\phi,\bar{\psi},\psi]}$ (15)
@@
$e^{-W_{\Lambda}^{\Lambda_{0}}[J]}=\int D\psi D\bar{\psi}D\phi e^{-\frac{1}{2}\int_{p}\phi(p)G_{B}^{-1}(\frac{p}{\Lambda})\phi(-p)-\int_{p}\bar{\psi}(p)G_{F}^{-1}(\frac{p}{\Lambda})\psi(-p)+\int_{p}J(p)\phi(-p)+\int_{p}\bar{\eta}(p)\psi(-p)+\int_{p}\eta(p)\bar{\psi}(-p)-S_{int}[\phi,\bar{\psi},\psi]}$ (16)
$=$ $\int D\psi D\bar{\psi}e^{-\int_{p}\bar{\psi}(p)G_{F}^{-1}(\frac{p}{\Lambda})\psi(-p)+\int_{p}\bar{\eta}(p)\psi(-p)+\int_{p}\eta(p)\bar{\psi}(-p)}$
×$\times e^{-\frac{1}{2}\int_{p}J(p)G_{B}(\frac{p}{\Lambda})J(-p)}e^{\frac{1}{2}\int_{p}G_{B}(\frac{p}{\Lambda})\frac{\delta}{\delta\phi(p)}\frac{\delta}{\delta\phi(-p)}e^{-S_{int}[\phi,\bar{\psi},\psi]}}\bigg{|}_{\phi=G_{B}(\frac{p}{\Lambda})J(p)}.$
@@
$e^{-W_{\Lambda}^{\Lambda_{0}}[J]}=\int D\psi D\bar{\psi}D\phi e^{-\frac{1}{2}\int_{p}\phi(p)G_{B}^{-1}(\frac{p}{\Lambda})\phi(-p)-\int_{p}\bar{\psi}(p)G_{F}^{-1}(\frac{p}{\Lambda})\psi(-p)+\int_{p}J(p)\phi(-p)+\int_{p}\bar{\eta}(p)\psi(-p)+\int_{p}\eta(p)\bar{\psi}(-p)-S_{int}[\phi,\bar{\psi},\psi]}$ (17)
$=$ $e^{-\frac{1}{2}\int_{p}J(p)G_{B}(\frac{p}{\Lambda})J(-p)}\bigg{[}e^{-\int_{p}\bar{\eta}(p)G_{F}\eta(-p)}e^{-\int_{p}G_{F}\frac{\delta}{\delta\bar{\psi}(p)}\frac{\delta}{\delta\psi(-p)}}e^{-\frac{1}{2}\int_{p}G_{B}(\frac{p}{\Lambda})\frac{\delta}{\delta\phi(p)}\frac{\delta}{\delta\phi(-p)}}e^{-S_{int}[\phi,\bar{\psi},\psi]}\bigg{]}\bigg{|}_{\phi=G_{B}J,\bar{\psi}=G_{F}\eta,\psi=G_{F}\bar{\eta}}.$
@@
$\frac{\partial}{\partial\Lambda}e^{-S_{int}[\phi,\bar{\psi},\psi,\Lambda]}=\frac{\partial}{\partial\Lambda}\big{(}e^{-\frac{1}{2}\int_{p}J(p)G_{B}(\frac{p}{\Lambda})J(-p)}e^{-\int_{p}\bar{\eta}(p)G_{F}(\frac{p}{\Lambda})\eta(-p)}e^{-\int_{p}G_{F}(\frac{p}{\Lambda})\frac{\delta}{\delta\bar{\psi}(p)}\frac{\delta}{\delta\psi(-p)}}e^{-\frac{1}{2}\int_{p}G_{B}(\frac{p}{\Lambda})\frac{\delta}{\delta\phi(p)}\frac{\delta}{\delta\phi(-p)}}e^{-S_{int}[\phi,\bar{\psi},\psi]}\big{)},$ (19)
@@
$-\Lambda\frac{\partial S_{int}}{\partial\Lambda}=$ $\int_{p}\Lambda\frac{\partial G_{F}(\frac{p}{\Lambda})}{\partial\Lambda}\bigg{(}\frac{\delta S_{int}}{\delta\bar{\psi}(p)}\frac{\delta S_{int}}{\delta\psi(-p)}-\frac{\delta^{2}S_{int}}{\delta\bar{\psi}(p)\delta\psi(-p)}\bigg{)}+\frac{1}{2}\int_{p}\Lambda\frac{\partial G_{B}(\frac{p}{\Lambda})}{\partial\Lambda}\bigg{(}\frac{\delta S_{int}}{\delta\phi(p)}\frac{\delta S_{int}}{\delta\phi(-p)}-\frac{\delta^{2}S_{int}}{\delta\phi(p)\delta\phi(-p)}\bigg{)}$ (20)
$=$ $\int_{p}G_{F}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S_{int}}{\delta\bar{\psi}(p)}\frac{\delta S_{int}}{\delta\psi(-p)}-\frac{\delta^{2}S_{int}}{\delta\bar{\psi}(p)\delta\psi(-p)}\bigg{)}+\frac{1}{2}\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S_{int}}{\delta\phi(p)}\frac{\delta S_{int}}{\delta\phi(-p)}-\frac{\delta^{2}S_{int}}{\delta\phi(p)\delta\phi(-p)}\bigg{)},$
@@
The first step of Wilson’s Renormalization Group program is now complete. The first step already performed above, involves integrating out the high momentum modes of the fields, which is known as coarse-graining or blocking, this generates an equivalent effective action with reduced physical cutoff scale. Now, this has to be followed by the second step which is rescaling of the theory’s action parameters (variables), by rescaling the momenta back to the original physical cutoff scale including appropriate rescaling of fields in order for the resulting equations to be cutoff scale independent, which also ensures the equations (theory) to have fixed points solutions relating to scale-invariant, massless renormalized theory. In the first step, the fields $\phi(p)$ are integrated over the range $e^{-t}<|p|\leq 1$ such that the partition function is left invariant, and the second step involves rescaling, a change by a factor of $e^{-t}$ in all linear dimensions to restore the original scale $\Lambda$ of the system, i.e $p\rightarrow p^{\prime}=e^{t}p$, where parameter $t$ is an infinitesimal parameter.
@@
and applying this form of rescaling on Polchinski’s ERG flow equation (20) for a case of single self-interacting Bosonic fields leads to the full rescaled Polchinski’s ERG equation,
$\big{(}\Lambda\partial_{\Lambda}+d_{\phi}\bigtriangleup_{\phi}+\bigtriangleup_{\partial_{\phi}}-d\big{)}S[\phi,\Lambda]=\frac{1}{2}\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S[\phi,\Lambda]}{\delta\phi(p)}\frac{\delta S[\phi,\Lambda]}{\delta\phi(-p)}-\frac{\delta^{2}S[\phi,\Lambda]}{\delta\phi(p)\delta\phi(-p)}\bigg{)}$ (24)
$\big{(}\partial_{t}+d_{\phi}\bigtriangleup_{\phi}+\bigtriangleup_{\partial_{\phi}}-d\big{)}S[\phi,t]=\frac{1}{2}\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S[\phi,t]}{\delta\phi(p)}\frac{\delta S[\phi,t]}{\delta\phi(-p)}-\frac{\delta^{2}S[\phi,t]}{\delta\phi(p)\delta\phi(-p)}\bigg{)},$
where $\partial_{t}=\Lambda\frac{\partial}{\partial\Lambda}$ is the time derivative, $\Delta_{\partial_{\phi}}=d+\int_{p}\phi_{p}p\cdot\partial_{p}\frac{\delta}{\delta\phi_{p}}$, $\Delta_{\phi}=\int_{p}\phi_{p}\frac{\delta}{\delta\phi_{p}}$, $d_{\phi}=\frac{1}{2}(d-2+\eta_{\phi})$ and $\eta_{\phi}$ is the anomalous dimension of the scalar field $\phi(p)$. The details regarding how the results in equation (24) were obtained are provided in the Appendix. Treating the Fermionic fields in similar fashion, in the case of self-interacting spinless Fermionic fields, now $S_{int}$ will be $S[\psi,\bar{\psi},\Lambda]$, and the final derivation results for the Polchinski’s ERG flow equation is,
@@
where $\Delta_{\psi\bar{\psi}}=\int_{p}\psi(p)\frac{\delta}{\delta\psi(-p)}+\int_{p}\bar{\psi}(p)\frac{\delta}{\delta\bar{\psi}(-p)}$, $\Delta_{\partial_{\psi\bar{\psi}}}=d+\int\frac{d^{d}p}{(2\pi)^{d}}\psi(p)p\cdot\frac{\partial}{\partial p}\frac{\delta}{\delta\psi(-p)}+\int\frac{d^{d}p}{(2\pi)^{d}}\bar{\psi}(p)p^{\cdot}\frac{\partial}{\partial p}\frac{\delta}{\delta\bar{\psi}(-p)}$, $d_{\psi\bar{\psi}}=\frac{1}{2}(d-1+\eta_{\psi})$ and $\eta_{\psi}$ is the anomalous dimensions of the Fermionic fields. Now, the general rescaled Polchinski’s ERG flow equation for the interactions between the single Bosonic and spinless Fermionic fields together will be,

### v2

@@ Abstract
By simply applying the Local Potential Approximation (LPA) on the Polchinski’s Exact Renormalization Group (ERG) flow equation for single Bosonic and spinless Fermionic fields, and initially considering only the coarse-graining (blocking) aspect of Wilson’s Renormalization Group program. Within the LPA limit the Polchinski’s ERG flow equation simplifies into a heat differential equation for both the Bosonic and Fermionic fields. Solving the differential equations leads to logarithmic interactions (logarithmic vertex function) in both the Bosonic and Fermionic fields at their fixed points.
preprint: APS/123-QED
I Introduction
Within the Framework of Exact Renormalization Group (ERG) methods [1, 2, 3, 4], with a particular focus on Polchinski’s Exact Renormalization Group Equation (ERGE) [4], which is the Wilson’s Exact Renormalization Group (ERG) [1] derived in the field theory context. We study a system described by a general effective action consisting of single Bosonic and spinless Fermionic field modes including their interactions. Working in Euclidean momentum space where the theory is regularized by employing the cutoff functions of generic form $K_{UV}(p,\Lambda)$ in both fields, whereby the cutoff functions suppress field modes with momentum $p$ that are above the physical UV (ultra-violet) cutoff scale $\Lambda$. The Polchinski’s ERGE (Exact Renormalization Group Equation) emanate from requiring that as the scale $\Lambda$ is lowered by integrating (course-graining) high momentum modes, the effective action (vertex functions) of the theory should be varied in such a way that the partition function of the theory remains unchanged. The variation of the effective action due to lowering of the physical cutoff scale is governed by the Renormalization Group (RG) equation, which is an exact differential equation subsequently giving rise to ERG methods. ERG methods in field theory are exact non-perturbative formulations of RG flow physics encoding it as a differential variation of physical scale and the effective action. They describe the evolution of the renormalized effective action along the RG flow trajectory in terms of differential equations. From the theory setup, these equations arise from the decimation (integration) of fields modes momenta, followed by lowering of the momentum cutoff and then rescaling the action parameters while enforcing the physical observables to remain invariant by requiring the partition function to remain invariant. Usually the RG flow has two physical cutoff scales, one is a high energy scale denoted as the UV cutoff scale, and the second one is a low energy scale denoted as an IR (infrared) cutoff scale, and the RG flow will normally be a flow from the UV (high energy) limit towards the IR (low energy) limit. A generic RG flow equation has the form
@@ I Introduction
Generally solving such an equation specially in the ERGE context can be challenging, since a handful of ERG’s problems end up as integro-differential equations [1, 2, 3, 4], with Polchinski’s ERG flow equation being in this category. Therefore finding an exact non-perturbative general solution from these kind of equations generally it is a difficult tasks, in practice one ends up with an infinite tower of equations and has to either truncate the equations by performing certain approximations. Several non-perturbative approximation solutions and methods are possible, such as derivative expansion, which involves expanding the equations in powers of momenta and this has shown to yield promising results for calculating critical exponents and renormalization group flows equations in $2\leq d\leq 4$ scalar fields theories [5, 6, 7, 8]. The simplest non-perturbative approximation from such power momentum expansion is obtained from taking the leading term (zeroth order term of the expansion), which is the Local Potential Approximation (LPA) expansion. Here we will make use of such power momentum expansion, restrict the computations within the LPA limit. As a start we avoid the consideration of next to leading order terms and other higher order terms in momentum or derivative expansion since these terms are known to yield physical solutions (physical observables) which are scheme dependent (cutoff scale dependent) [8, 9, 10], which makes it difficult to compute physical observables such as the critical exponents ambiguously. Whilst the advantage with the leading order solutions is that, the LPA physical observables solutions are scheme independent and can be computed unambiguously.
It is known and alluded [11] that the Polchinski’s ERG flow equation can be written as a functional heat equation. However, presently it appears there is lack of visible available work towards unpacking this alluded notion at a basic level. Clearly the notion that the Polchinski’s ERG flow equation can be represented as a functional heat equation hasn’t received much attention, and consequently not much work has been done to unpack and elucidate this notion with simple theoretical approaches. Therefore this paper aims to address this, by providing a simple connection or relation between the Polchinski’s ERG flow equation and the well known standard heat equation within the LPA limit. Utilizing this notion, in one of our main results in section 3 at equation (44) in the case of single Bosonic fields, we show that at low energy after relaxing the rescaling steps of Wilson’s Renormalization Group program and only implement the coarse-graining step, subsequently the Polchinski’s ERG flow equation within the LPA limit simplifies into a standard heat equation,
@@
where $U(z,t)=e^{-V(z,t)}$, $t$ is a dimensionless parameter, $z$ represent Bosonic fields and $V(z,t)$ is the vertex or interaction function. We will explain more about the results and how they are obtained in later sections of the paper. It is worth mentioning that the results are interesting due to their simplistic nature of relating the Polchinski’s ERG flow equation within the LPA limit to a heat equation, this can potentially open a new front to generally study ERG equations through the available technology and language already well developed and established in studying heat equations. Subsequently the solution to this standard heat equation at the fixed point, which is related to the steady state solution in the language of heat equations, we find that the fixed point solution is governed by logarithmic interactions,
@@
where $V_{*}(z,t)$ is the interaction function at the fixed point obtained at $t=\frac{z^{2}}{2\lambda}$. The solution is interesting since fixed points normally signals a Conformal Field Theory (CFT), therefore at this low energy limit we have established a logarithmic interaction CFT for the single Bosonic fields. Contrary to perturbing the solution in the vicinity of the Gaussian fixed point [12, 13, 14, 5], here we perturb the solution in the vicinity of the logarithmic fixed point obtained above. For the perturbation recipe we use a small parameter to linearize the Polchinski’s ERG flow equation in the vicinity of the logarithmic fixed point, but here we implement both rescaling and coarse-graining steps on the Polchinski’s ERG flow equation in order to find non-trivial fixed-point solutions. This leads to an infinite set tower of Hermite differentials at each expansion order in parameter $t$, with the Hermite polynomials being solutions to these differential equations. Treating the Fermionic fields in similar fashion, in section 4 at equation (82) we obtain the following analogue of a heat equation in Grassmann variables,
@@
The paper is organised as follows. In section 2 we discuss the derivation of Polchinski’s ERG flow equation for a system containing both the single Bosonic and spinless Fermionic fields, and only considering the coarse-graining step of Wislon’s Renormalization Group program. In section 3 we additionally consider the rescaling step of Wilson’s Renormalization Group program on the Polschinki’s ERG flow equation, we apply the rescaling step and take the LPA limit of the Polchinski’s ERG flow equation, then the differential equation emanating from Polchinki’s ERG flow equation at the LPA limit is solved, followed by finding the eigenvalues, eigenfunctions and the critical exponents for the Bosonic fields sector. Then section 4 deals with solving the Polchinski’s ERG flow equation within the LPA limit in the Fermionic fields sector.
II Derivation of Polchinski’s ERGE for Bosonic and Fermionic Fields.
@@
$-\Lambda\frac{\partial S_{int}[\phi,\Lambda]}{\partial\Lambda}=\frac{1}{2}\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta^{2}S_{int}[\phi,\Lambda]}{\delta\phi(p)\delta\phi(-p)}-\frac{\delta S_{int}[\phi,\Lambda]}{\delta\phi(p)}\frac{\delta S_{int}[\phi,\Lambda]}{\delta\phi(-p)}\bigg{)},$ (14)
@@
$e^{-W_{\Lambda}^{\Lambda_{0}}[J]}=\int D\psi D\bar{\psi}D\phi e^{-\frac{1}{2}\int_{p}\phi(p)G_{B}^{-1}(\frac{p}{\Lambda})\phi(-p)}$ (15)
×$\times e^{-\int_{p}\bar{\psi}(p)G_{F}^{-1}(\frac{p}{\Lambda})\psi(-p)+\int_{p}J(p)\phi(-p)+\int_{p}\bar{\eta}(p)\psi(-p)+\int_{p}\eta(p)\bar{\psi}(-p)}$
×$\times e^{-S_{int}^{\Lambda_{0}}[\phi,\bar{\psi},\psi]}$
@@
$e^{-W_{\Lambda}^{\Lambda_{0}}[J]}$ $=\int D\psi D\bar{\psi}e^{-\int_{p}\bar{\psi}(p)G_{F}^{-1}(\frac{p}{\Lambda})\psi(-p)+\int_{p}\bar{\eta}(p)\psi(-p)}$ (16)
×$\times e^{\int_{p}\eta(p)\bar{\psi}(-p)-\frac{1}{2}\int_{p}J(p)G_{B}(\frac{p}{\Lambda})J(-p)}$
@@
$e^{-W_{\Lambda}^{\Lambda_{0}}[J]}=e^{-\frac{1}{2}\int_{p}J(p)G_{B}(\frac{p}{\Lambda})J(-p)}$ (17)
×$\times\bigg{[}e^{-\int_{p}\bar{\eta}(p)G_{F}\eta(-p)}e^{-\int_{p}G_{F}\frac{\delta}{\delta\bar{\psi}(p)}\frac{\delta}{\delta\psi(-p)}}$
×$\times e^{-\frac{1}{2}\int_{p}G_{B}(\frac{p}{\Lambda})\frac{\delta}{\delta\phi(p)}\frac{\delta}{\delta\phi(-p)}}e^{-S_{int}[\phi,\bar{\psi},\psi]}\bigg{]}\bigg{|}_{\phi=G_{B}J,\bar{\psi}=G_{F}\eta,\psi=G_{F}\bar{\eta}}.$
@@
$e^{-S_{int}[\phi,\bar{\psi},\psi,\Lambda]}=e^{-\frac{1}{2}\int_{p}J(p)G_{B}(\frac{p}{\Lambda})J(-p)}e^{-\int_{p}\bar{\eta}(p)G_{F}(\frac{p}{\Lambda})\eta(-p)}$ (18)
×$\times e^{-\int_{p}G_{F}(\frac{p}{\Lambda})\frac{\delta}{\delta\bar{\psi}(p)}\frac{\delta}{\delta\psi(-p)}}e^{-\frac{1}{2}\int_{p}G_{B}(\frac{p}{\Lambda})\frac{\delta}{\delta\phi(p)}\frac{\delta}{\delta\phi(-p)}}e^{-S_{int}[\phi,\bar{\psi},\psi]},$
@@
$\frac{\partial}{\partial\Lambda}e^{-S_{int}[\phi,\bar{\psi},\psi,\Lambda]}$ $=\frac{\partial}{\partial\Lambda}\big{(}e^{-\frac{1}{2}\int_{p}J(p)G_{B}(\frac{p}{\Lambda})J(-p)}$ (19)
×$\times e^{-\int_{p}\bar{\eta}(p)G_{F}(\frac{p}{\Lambda})\eta(-p)}e^{-\int_{p}G_{F}(\frac{p}{\Lambda})\frac{\delta}{\delta\bar{\psi}(p)}\frac{\delta}{\delta\psi(-p)}}$
×$\times e^{-\frac{1}{2}\int_{p}G_{B}(\frac{p}{\Lambda})\frac{\delta}{\delta\phi(p)}\frac{\delta}{\delta\phi(-p)}}e^{-S_{int}[\phi,\bar{\psi},\psi]}\big{)},$
@@
$-\Lambda\frac{\partial S_{int}}{\partial\Lambda}=$ $\int_{p}G_{F}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S_{int}}{\delta\bar{\psi}(p)}\frac{\delta S_{int}}{\delta\psi(-p)}-\frac{\delta^{2}S_{int}}{\delta\bar{\psi}(p)\delta\psi(-p)}\bigg{)}+\frac{1}{2}\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S_{int}}{\delta\phi(p)}\frac{\delta S_{int}}{\delta\phi(-p)}-\frac{\delta^{2}S_{int}}{\delta\phi(p)\delta\phi(-p)}\bigg{)},$ (20)
@@
The first step of Wilson’s Renormalization Group program is now complete. The first step already performed above, involves integrating out the high momentum modes of the fields, which is known as coarse-graining or blocking, this generates an equivalent effective action with reduced physical cutoff scale. Now, this has to be followed by the second step which is rescaling of the theory’s action parameters (variables), by rescaling the momenta back to the original physical cutoff scale including appropriate rescaling of fields in order for the resulting equations to be cutoff scale independent, which also ensures the equations (theory) to have fixed points solutions relating to scale-invariant, massless renormalized theory. In the first step, the fields $\phi(p)$ are integrated over the range $e^{-t}<|p|\leq 1$ such that the partition function is left invariant, and the second step involves rescaling, a change by a factor of $e^{-t}$ in all linear dimensions to restore the original scale $\Lambda$ of the system, i.e $p\rightarrow p^{\prime}=e^{t}p$, where the parameter $t$ is an infinitesimal parameter.
@@
$S[\phi,\Lambda]=$ $\sum_{m}\int\prod_{i=1}^{m}\frac{d^{d}p_{i}}{(2\pi)^{d}}\mathcal{L}_{m,0}(p_{1},\cdots,p_{m},\Lambda)$ (22)
×$\times\phi(p_{1})\cdots\phi(p_{m})\delta^{d}(p_{1}+\cdots+p_{m}).$
@@
and applying this form of rescaling on Polchinski’s ERG flow equation (20) for a case of single self-interacting Bosonic fields leads to the full rescaled Polchinski’s ERG flow equation,
$\big{(}\partial_{t}+d_{\phi}\bigtriangleup_{\phi}+\bigtriangleup_{\partial_{\phi}}-d\big{)}S[\phi,t]=\frac{1}{2}\int_{p}G_{B}^{\prime}\big{(}\frac{p}{\Lambda}\big{)}\bigg{(}\frac{\delta S[\phi,t]}{\delta\phi(p)}\frac{\delta S[\phi,t]}{\delta\phi(-p)}-\frac{\delta^{2}S[\phi,t]}{\delta\phi(p)\delta\phi(-p)}\bigg{)},$ (24)
where $\partial_{t}=\Lambda\frac{\partial}{\partial\Lambda}$ is the time derivative,
$\Delta_{\partial_{\phi}}=d+\int_{p}\phi_{p}p\cdot\partial_{p}\frac{\delta}{\delta\phi_{p}},$ (25)
$\Delta_{\phi}=\int_{p}\phi_{p}\frac{\delta}{\delta\phi_{p}},$
$d_{\phi}=\frac{1}{2}(d-2+\eta_{\phi})$
and $\eta_{\phi}$ is the anomalous dimension of the scalar field $\phi(p)$. The details regarding how the results in equation (24) were obtained are provided in the Appendix. Treating the Fermionic fields in similar fashion, in the case of self-interacting spinless Fermionic fields, now $S_{int}$ will be $S[\psi,\bar{\psi},\Lambda]$, and the final derivation results for the Polchinski’s ERG flow equation is,
@@
where
$\Delta_{\partial_{\psi\bar{\psi}}}=d+\int\frac{d^{d}p}{(2\pi)^{d}}\psi(p)p\cdot\frac{\partial}{\partial p}\frac{\delta}{\delta\psi(-p)}$ (27)
$\quad+\int\frac{d^{d}p}{(2\pi)^{d}}\bar{\psi}(p)p^{\cdot}\frac{\partial}{\partial p}\frac{\delta}{\delta\bar{\psi}(-p)}$
$\Delta_{\psi\bar{\psi}}=\int_{p}\psi(p)\frac{\delta}{\delta\psi(-p)}+\int_{p}\bar{\psi}(p)\frac{\delta}{\delta\bar{\psi}(-p)}$ (28)
$d_{\psi\bar{\psi}}=\frac{1}{2}(d-1+\eta_{\psi})$
and $\eta_{\psi}$ is the anomalous dimensions of the Fermionic fields. Now, the general rescaled Polchinski’s ERG flow equation for the interactions between the single Bosonic and spinless Fermionic fields together will be,